OpenStudy (anonymous):

what is the prime factorization of 225? a.2x3x5x5 b.9x25 c.3x3x5x5

OpenStudy (anonymous):

option c

OpenStudy (ash2326):

\[225=9*25=3^2*5^2 \] or \[3*3*5*5\] option C

OpenStudy (anonymous):

yes C

OpenStudy (anonymous):

its "c"